OpenStudy (scarlettfarra2000):

please help me 1) - - to convert 0.92 to fraction, Lauren wrote n=0.92 as her first step for her second step what should she multiply both sides of the equation 10 100 1000 10000 2) which number is between 1/7 and 0.2? 9/35 6/35 2/9 1/9 3) - which number is between 6/11 and 0.67 11/18 69/99 24/33 26/55 4) what is 3/13 as a decimal? ---- 0.230769 -- 0.230769 0.230769 ------- 0.230769

OpenStudy (scarlettfarra2000):

5) which fraction shows 2.8 as quotient of two intergers? 14/5 17/5 9/4 9/2

OpenStudy (civicsiscool44):

number two is D

OpenStudy (mustbethestars):

This is the wrong section ;-;. Number 1 is B because a percent goes up too 100. So 1/100 would be 1% The easiest way to do numbber 2 is to turn 1/7 into a fraction. So in your calculator do 1 divided by 7. That would be 0.14. Now turn all the answers into fractions too. We are looking for a number between 0.14 and 0.20. 9/35= 0.26. We can cross A off the list. 6/35 = 0.17 so we can leave that on the list. 9/4 = 2.25 so we can cross C off the list. 9/2 = 4.5 . So the Only possible answer is B

OpenStudy (mustbethestars):

Well anyways we do Number 3 the same way we did Number 2. The decimal form of 6/11 is 0.55. So a number between 0.55 and 0.67. 11/18= 0.61. So A is a possible answer. 69/99 = 0.70. So B can be crossed off the list. 24/33 is 0.73 so C can be crossed off the list. 26/55 = 0.47. So D can be crossed off the list. Leaving A the only possible answer. For number 4 all answer choices are the same and correct, I have never seen that before though. But when you have a fraction that needs to be a decimal it just becomes regular division. For number 5, just divide all answer choices and see which one is 2.8. 14/5 = 2.8. 17/5 = 3.4. 9/4 = 2.25. 9/2 = 4.5. So the answer is A.

OpenStudy (wwhitlock):

@ScarlettFarra2000 You got a bunch of help from @mustbethestars. And you gave him nothin'. Give a medal or something.

OpenStudy (scarlettfarra2000):

Thank you @mustbethestars for your help sorry I didn't say that before I stepped away for a second

OpenStudy (mustbethestars):

It's okay. No problem. If you need any more help message or tag me :)